OpenStudy (anonymous):

Find all the real square roots of 0.0081

OpenStudy (helder_edwin):

every non-negative real number has only ONE square root

OpenStudy (zpupster):

sqrt 81 = 9 can you go from there

OpenStudy (anonymous):

so .09 and -.09

OpenStudy (helder_edwin):

good point \[\large 0,0081=\frac{81}{10000} \]

OpenStudy (helder_edwin):

so \[\large \sqrt{0,0081}=\sqrt{\frac{81}{10000}} \]

OpenStudy (mathmale):

Edwin, that's a very helpful approach.

OpenStudy (zpupster):

so .09 and -.09 is right